OpenStudy (anonymous):

Hi, I really need help UNDERSTANDING this question!What is the final velocity of a car that accelerated 10 m/s2 from rest and traveled 180m? I mean, whats the formula that you do it and and the steps you take? thank you very much

OpenStudy (anonymous):

The answer is 60, but why?

OpenStudy (paxpolaris):

http://en.wikipedia.org/wiki/Equations_of_motion#Constant_linear_acceleration check out the sub-section: SUVAT equations I learned and used equations 1,2,& 4 in school.... eq. 4 is applicable in your case \[\Large v^2=u^2+2as\]

OpenStudy (anonymous):

\[v ^{2}=u ^{2}+2as\] initial velocity is 0 m/s, i.e u=0m/s \[v ^{2}=0+2as\] \[v ^{2}=2as\] and acceleration is 10 m/s^2 and distance traveled is 180 m, i.e a=10 m/s^2 and s=180 m \[v ^{2}=2*10*180\] \[v ^{2}=3600\] \[v =+60\] so final velocity is 60m/s (^_^)

OpenStudy (anonymous):

u can solve this in a quite simple way with avarge speed or with a graph.  the distance is the area under the line (simple triangle) that is \[V(final)*t/2=180 and V(final)=at \] solve it for t and u r almost done.. it is just the same using the avarge speed